OpenStudy (abymartinez15):

Please help............................ Danika concludes that the following functions are inverses of each other because f(g(x)) = x. Do you agree with Danika? Explain your reasoning. f(x) = |x| g(x) = –x

OpenStudy (welshfella):

replace the x in f(x) by the expression for g(x). What do you get

OpenStudy (abymartinez15):

so i would have to solve for g(x)?

OpenStudy (welshfella):

Just replace the x in f(x) by g(x) that will give you an expression for f{(g)x)

OpenStudy (welshfella):

To give you a simple example if f(x) = x -1 and g(x) = 2x then f(g(x)) = 2x - 1

OpenStudy (welshfella):



OpenStudy (abymartinez15):

so x will be one right?

OpenStudy (welshfella):

No x is just a variable and can be any real value

OpenStudy (abymartinez15):

so it could be any number then

OpenStudy (welshfella):

yes but we are not solving an equation here we want to see if f(g(x)) for the given f(x) and g(x) equals x.

OpenStudy (abymartinez15):

okay, so f(g(x)) will have to give me a number?

OpenStudy (welshfella):

try doing it this way f(x) = |x| let y = f(x) so y = |x| Now solve for x

OpenStudy (abymartinez15):

that this equation will give me a number right?

OpenStudy (welshfella):

no it will give you an expression in x

OpenStudy (welshfella):

* in y

OpenStudy (abymartinez15):

i dont get it :(

OpenStudy (welshfella):

sorry but i have to go right now I'm sure youll get help from another person

OpenStudy (abymartinez15):

thank you anyway really appreciate it

OpenStudy (deaconvex11511):

derp